A. Find the length of the latus rectum of 4y = x2 + 4x + 8.

B. Find the radius of the circle with equation 4x2 + 4y2 - 17x + 17y = 19
Answer #1

A. Ginen equation 44 = x2 + 4x +8 4y = (x3+4x + 47+4 4y = (x + 2)2+4 (x+2)2 = 464-1) compare with x²= day Hence latus rectum